# LiD-FL: Towards List-Decodable Federated Learning

### v1

@@ B.3 Model Architecture and Hyperparameters
To introduce the structure of model, we adopt the compact notation in (Mhamdi, Guerraoui, and Rouault 2021):
@@
For the FEMNIST dataset, we use the stochastic gradient descent optimizer (SGD), with a learning rate of $0.01$ and a momentum coefficient of $0.9$. The number of local training rounds is $25$, and the batch size is $32$ for LR and $64$ for CNN, respectively. We use the minimal size of the global model list, i.e. $q=\lfloor 1/\gamma\rfloor$. The number of global training rounds is $1500$. For the CIFAR-10 dataset, the hyperparameters are the same as the CNN model on FEMNIST, except that the number of local training rounds is $20$ with a batch size of $32$, and the number of global training rounds is $2000$.
@@ B.4 Full Experimental Results
In this section we display the full experimental results, completing Section 5. We conduct experiments on three pairs dataset-model: FEMNIST-LR, FEMNIST-CNN, CIFAR-10-CNN. On FEMNIST, we simulate $3$ levels of Byzantine ratio to evaluate the performance of LiD-FL. For CIFAR-10, we choose the level $0.6$ to validate the effectiveness of LiD-FL. We use the test accuracy at the last training round as the evaluation criteria. The full results are presented in Table 3, Table 4 and Table 5, with each cell records the average and standard deviation of testing accuracy across $5$ runs.
@@
Rate Method Empire Gauss LF LIE OMN SF Wst
@@
Method Empire Gauss LF LIE OMN SF Wst
@@ B.5 Stability of LiD-FL
To demonstrate the influence of list size $q$ on the performance of LiD-FL, we conduct a grid search with $q$ in $[2,3,4,5]$. The experimental setup is the same as in Section 5.1, at a Byzantine fraction of $0.6$. For simplicity, we choose one representative attack for data poisoning and another for model poisoning. The results are illustrated in Table 6 and Figure 4. The performance of LiD-FL is relatively stable with respect to the list size on both models. Furthermore, larger list sizes result in reduced deviation across the test accuracy of 5 random experiments, indicating better stability.
@@ B.6 LiD-FL with Aggregators
We incorporate an aggregator into Algorithm 1, aiming to improve data parallelism of LiD-FL. The aggregator $h$ is a simple scheme for mean estimation, which takes $m$ high-dimensional vectors as input and outputs a vector with the same shape. Additionally, The aggregator needs to know the number of non-Byzantine clients $k$. We design two aggregation rules, the random $k$ nearest neighbors aggregation (RKNN) and the minimum enclosing ball aggregation (MEB), and conduct a preliminary experiment to evaluate the performance of LiD-FL with aggregators.
@@
We conduct experiments on FEMNIST using the same hyperparameter settings as described in Section B.3. The Byzantine fraction is $0.6$ and the voting attack is Worst. The performance comparison between naive LiD-FL and LiD-FL with aggregators is displayed in Table 7 and Figure 5. Table 7 punctuates the effectiveness of aggregators, with which the test accuracy of LiD-FL is improved overall. For the LR global model, both aggregators enhance the performance of LiD-FL. For the CNN global model, LiD-FL with both aggregators outperforms naive LiD-FL against Empire, Guass, LF and SF attacks. However, the performance of LiD-FL with MEB is inferior in the face of LIE and OMN attacks.In addition, LiD-FL with aggregators often shows better stability, supported by smaller standard deviation of test accuracy than naive LiD-FL. These phenomena require further in-depth study. We leave the problem of finding optimal aggregation rules for Lid-FL open.
@@
Model Method Empire Gauss LF LIE Omn SF Worst
LR LiD-FL 0.57 ± 0.03 0.55 ± 0.03 0.56 ± 0.03 0.57 ± 0.03 0.57 ± 0.03 0.56 ± 0.03 0.55
@@
CNN LiD-FL 0.72 ± 0.05 0.71 ± 0.04 0.72 ± 0.04 0.72 ± 0.03 0.73 ± 0.03 0.73 ± 0.03 0.71
@@
Table 7: Performance comparison of naive LiD-FL and LiD-FL plus aggregator on FEMNIST at a Byzantine fraction of $0.6$.
@@
3:     Select a starting point at random: $u_{i}\in A$.
4:     Find $N(u_{i})$, the set of $k$ nearest points in $A$ to $u_{i}$, breaking ties arbitrarily.
5:     $\mathcal{M}\leftarrow\mathcal{M}\cup\{\frac{1}{|N(u_{i})|}\sum_{u\in N(u_{i})}{u}\}$.
6:     $A\leftarrow A\backslash N(u_{i})$.
@@
Input: a set of $m$ vectors $U=\{u_{1},\dots,u_{m}\}$ in $\mathbb{R}^{d}$, and the number of normal points $k,k\leq m$.
@@
4:     for $i={1,\dots,|A|}$ do
5:        Find $N(u_{i})$, the set of $k$ nearest points in $A$ to $u_{i}$, breaking ties arbitrarily.
6:        $d=\max_{u\in N(u_{i})}{||u-u_{i}||}_{2}$.
7:        $R\leftarrow R\cup{d}$.
8:     end for
9:     $j=\operatorname*{arg\,min}_{i}R$.
10:     $\mathcal{M}\leftarrow\mathcal{M}\cup\{\frac{1}{|N(u_{j})|}\sum_{u\in N(u_{j})}{u}\}$.
11:     $A\leftarrow A\backslash N(u_{j})$.
12:  end while
13:  return  One aggregated update $\hat{o}$ chosen randomly from $\mathcal{M}$.

### v3

@@ B.3 Model Architecture and Hyperparameters
To introduce the structure of model, we adopt the compact notation in (Mhamdi et al. 2021):
@@
For the FEMNIST dataset, we use the stochastic gradient descent optimizer (SGD), with a learning rate of $0.01$ and a momentum coefficient of $0.9$. The initial momentum of local model is set to 00. The number of local training rounds is $25$, and the batch size is $32$ for LR and $64$ for CNN, respectively. We use the minimal size of the global model list, i.e. $q=\lfloor 1/\gamma\rfloor$. The number of global training rounds is $1500$. For the CIFAR-10 dataset, the hyperparameters are the same as the CNN model on FEMNIST, except that the number of local training rounds is $20$ with a batch size of $32$, and the number of global training rounds is $2000$.
@@ B.4 Full Experimental Results
In this section we display the full experimental results, completing Section 5 of the main paper. We conduct experiments on three pairs dataset-model: FEMNIST-LR, FEMNIST-CNN, CIFAR-10-CNN. On FEMNIST, we simulate $3$ levels of Byzantine ratio to evaluate the performance of LiD-FL. For CIFAR-10, we choose the level $0.6$ to validate the effectiveness of LiD-FL. We use the test accuracy at the last training round as the evaluation criteria. The full results are presented in Table 4, Table 5 and Table 6, with each cell records the average and standard deviation of testing accuracy across $5$ runs.
@@
Rate Method Empire Gauss LF LIE OMN SF Worst
@@
Method Empire Gauss LF LIE OMN SF Worst
@@ B.5 Stability of LiD-FL
To demonstrate the influence of list size $q$ on the performance of LiD-FL, we conduct a grid search with $q$ in $[2,3,4,5]$. The experimental setup is the same as in Section B.3, at a Byzantine fraction of $0.6$. For simplicity, we choose one representative attack for data poisoning and another for model poisoning. The results are illustrated in Table 7 and Figure 4. The performance of LiD-FL is relatively stable with respect to the list size on both models. Furthermore, larger list sizes result in smaller standard deviations across the test accuracy of $5$ runs, indicating better stability.
@@ B.6 LiD-FL with Aggregators
We incorporate an aggregator into LiD-FL, aiming to improve its data parallelism. The aggregator $h$ is a simple scheme for mean estimation, which takes $m$ high-dimensional vectors as input and outputs a vector with the same shape. Additionally, The aggregator needs to know the number of non-Byzantine clients $k$. We design two aggregation rules, the random $k$ nearest neighbors aggregation (RKNN) and the minimum enclosing ball aggregation (MEB), and conduct a preliminary experiment to evaluate the performance of LiD-FL with aggregators.
@@
We conduct experiments on FEMNIST using the same hyperparameter settings as described in Section B.3. The Byzantine fraction is $0.6$ and the voting attack is Worst. The performance comparison between naive LiD-FL and LiD-FL with aggregators is displayed in Table 8, Table 9 and Figure 5. Table 8 and Table 9 punctuate the effectiveness of aggregators, with which the test accuracy of LiD-FL is improved overall. For the LR global model, both aggregators enhance the performance of LiD-FL. For the CNN global model, LiD-FL with both aggregators outperforms naive LiD-FL against Empire, Guass, LF and SF attacks. However, the performance of LiD-FL with MEB is inferior in the face of LIE and OMN attacks.In addition, LiD-FL with aggregators often shows better stability, supported by smaller standard deviation of test accuracy than naive LiD-FL. These phenomena require further in-depth study. We leave the problem of finding optimal aggregation rules for LiD-FL open.
@@
Method Empire Gauss LF LIE Omn SF Worst
@@
Table 8: Performance comparison between naive LiD-FL and LiD-FL plus aggregator on FEMNIST with the LR model.
Method Empire Gauss LF LIE Omn SF Worst
@@
Table 9: Performance comparison between naive LiD-FL and LiD-FL plus aggregator on FEMNIST with the CNN model.
@@
3:     Select a starting point at random: $a_{i}\in A$.
4:     Let $N(a_{i})$ be the set of $k$ nearest points in $A$ to $a_{i}$, breaking ties arbitrarily.
5:     $\mathcal{M}\leftarrow\mathcal{M}\cup\{\frac{1}{|N(a_{i})|}\sum_{u\in N(a_{i})}{u}\}$.
6:     $A\leftarrow A\backslash N(a_{i})$.
@@
Input: a set of $m$ vectors $U=\{u_{1},u_{2},\dots,u_{m}\}$ in $\mathbb{R}^{d}$, and the number of normal points $k,k\leq m$.
@@
4:     for $i\in 1,2,\dots,|A|$ do
5:        Let $N(a_{i})$ be the set of $k$ nearest points in $A$ to $a_{i}$, breaking ties arbitrarily.
6:        $r_{i}\leftarrow\max_{u\in N(a_{i})}{||u-a_{i}||}_{2}$.
7:     end for
8:     $j\leftarrow\mathop{\arg\min}\limits_{i}\{r_{i}\}_{i=1}^{|A|}$.
9:     $\mathcal{M}\leftarrow\mathcal{M}\cup\{\frac{1}{|N(a_{j})|}\sum_{u\in N(a_{j})}{u}\}$.
10:     $A\leftarrow A\backslash N(a_{j})$.
11:  end while
12:  return  One aggregated update $\hat{o}$ chosen randomly from $\mathcal{M}$.